Your old toy train from pre-school days consists o
Answer #1

as all the cars are attached so the they have comman acceleration ( lets say a)

as Fnet = ma

5 - fr = (m1+m2+m3) a      

{ fr = friction force = mu * mg = 0.14 * ( 0.4+0.4+0.1) *9.81 = 1.23 N}

5 - 1.23 = ( 0.4+0.4+0.1) a

a = 4.18 m/s2

1) now take m1 only

F - FTA - Fr1 = m1 a

{ Fr1 = mu * 0.4 *9.81 =0.55 N}

5 - FTA- 0.55 = 0.4 * 4.18

FTA = 2.78 N                 ..................(1)

2) now on m3

FTB - Fr = m3 *a

{ Fr = mu m3 g = 0.14 * 0.1 *9.81 = 0.137 N}

so FTB -0.137 = 0.1 * 4.18

FTB = 0.55 N

so    FTA - FTB = 2.78 -0.55

= 2.22 N answer